# Zehan Li, Genliang Zhao, Jian Kang, Yongxiang Li, Jie Li

### v1

@@ 2.1 Modality-Alignment Training
Inspired by previous work on modality alignment [15; 16], our framework adopts speech-text continuation pairs as the core training data for speech-text alignment. In this stage, the LLM takes speech-derived query embeddings as input and generates text continuations that are generated by the LLM when it consumes the corresponding speech transcripts. Two complementary strategies are proposed to construct a sufficiently large and diverse dataset of such pairs. We first exploit transcripts from large-scale automatic speech recognition (ASR) corpora, which encompass Mandarin, English, and multiple dialects. These transcripts are extended by prompting the LLM to generate contextually coherent continuation text. This approach leverages the acoustic-semantic correlations embedded in ASR corpora. Second, we employ TTS to convert semantically coherent sentences into speech audio, while the LLM simultaneously generates continuation text for these sentences. These two processes ensure that the training data maintains both broad semantic diversity and syntactic coherence. To encode modality-specific characteristics, we prepend natural language descriptors (e.g., dialect specifications, emotional cues) to the input text during continuation generation. This explicit contextual guidance enables the model to learn modality-dependent representations for dialectal variations and affective expressions. This structured approach addresses the challenge of data scarcity in cross-modal alignment.
@@ 2.2 Speech-Generation Training
For parameter optimization, we adopt a modular training strategy where the speech encoder, projector, and foundational layers of the LLM are frozen to preserve pre-trained acoustic and linguistic representations. Meanwhile, the top-k semantic-to-speech token mapping layers are fine-tuned to adapt cross-modal interactions. A critical innovation is the multi-token prediction mechanism, which propagates temporal coherence by concatenating the current-step output token embeddings with next-step input features within the frozen upper network parameters. This fusion of sequential contextual information generates a temporally coherent latent representation, which guides the iterative generation of subsequent speech tokens.

### v2

# Speech Generation via A Dual-Branch LLM

@@ 2.1 Modality-Alignment Training
Inspired by previous work on modality alignment [15; 16], our framework adopts speech-text continuation pairs as the core training data for speech-text alignment. In this stage, the LLM takes speech-derived query embeddings as input and generates text continuations that are generated by the LLM when it consumes the corresponding speech transcripts. Two complementary strategies are proposed to construct a sufficiently large and diverse dataset of such pairs. We first exploit transcripts from large-scale automatic speech recognition (ASR) corpora, which encompass Mandarin, English, and multiple dialects. These transcripts are extended by prompting the LLM to generate contextually coherent continuation text. This approach leverages the acoustic-semantic correlations embedded in ASR corpora. Second, we employ TTS to convert semantically coherent sentences into speech audio, while the LLM simultaneously generates continuation text for these sentences. These two processes ensure that the training data maintains both broad acoustic and semantic diversity. To encode modality-specific characteristics, we prepend natural language descriptors (e.g., dialect specifications, emotional cues) to the input text during continuation generation. This explicit contextual guidance enables the model to learn modality-dependent representations for dialectal variations and affective expressions. This structured approach addresses the challenge of data scarcity in cross-modal alignment.
@@ 2.2 Speech-Generation Training
For parameter optimization, we adopt a modular training strategy where the speech encoder, projector, and foundational layers of the LLM are frozen to preserve pre-trained acoustic and linguistic representations. Meanwhile, the top-k semantic-to-speech token mapping layers are fine-tuned to adapt cross-modal interactions. In our implementation, we set $N=\lfloor M/2\rfloor$ and $K=M-N$, where $M$ is the total number of the LLM layers. We resort the further investigation of other possible settings to future work. Furthermore, a critical innovation is the multi-token prediction mechanism, which propagates temporal coherence by concatenating the current-step output token embeddings with next-step input features within the frozen upper network parameters. This fusion of sequential contextual information generates a temporally coherent latent representation, which guides the iterative generation of subsequent speech tokens.